A train runs in uniform speed and cover 360Km. If speed reduce by 8Km/hour, than it will take 4hours more than ptevious time to cover 360km. find the speed of train.

Answer by mananth(16946)

distance = 360 km
speed = x km/h
time taken = 360/x
..
speed reduced by 8
so speed = x-8 km/h
..
time = 360/x-8
reduced speed time - original time = 4 hours
..
360/(x-8)-360/x = 4
LCD = x(x-8)
360x-360(x-8)/(x(x-8))=4
360x-360x+2880= 4x(x-8)
2880=4x^2-32x
4x^2-32x -2880=0
/4
x^2-8x-720=0
use quadratic equation formula
discriminant = 2944
x1=((8+sqrt(2944))/2
x1=31.12 mph the original speed
x2=((8-sqrt(2944))/2
x2 is negative so ignore.
m.ananth@hotmail.ca